You are managing a pension fund with a value of $300 million and a beta of 1.07. You are concerned about a market decline and wish to hedge the portfolio. You have decided to use SPX calls. How many contracts do you need if the delta of the call option is 0.62 and the S&P Index is currently at 2030?
Answer #1

Each contract is for 100 shares of the underlying.

Notional value of the S&P index = 2030 * 100 = $203,000

Number of call contracts = (Pention fund value * Beta of the portfolio)/(Notional value of the index * call option delta)

Number of call contracts = (300,000,000 * 1.07)/(203,000 * 0.62)

Number of call contracts = 2,550.452884157

Can you please upvote? Thank You :-)